ELECTRONIC AND PHOTOELECTRON SPECTROSCOPY OF TOLUENE

\underline{ADRIAN M. GARDNER}, ALISTAIR M. GREEN, VICTOR TAMÉ-REYES and TIMO-THY G. WRIGHT, School of Chemistry, University of Nottingham, University Park, Nottingham, NG7 2RD, UK.

Electronic and photoelectron spectra of toluene are presented and discussed. The utilization of a recently reported scheme for assigning the normal vibrations of substituted benzenes allows these spectra to be compared to those of other molecules with unprecedented clarity.^{a} Changes in vibrational activity within a series of substituted benzene molecules will be discussed, specifically the increased rate of intramolecular vibrational energy redistribution observed in molecules where the substituent is a methyl group.

aA. M. Gardner and T. G. Wright, J. Chem. Phys., 135, 114305 (2011)